OpenStudy (anonymous):

tom quig traveled 280 miles east of saint louis. for most of the trip he averaged 60mph, but for one period of time he was slowed to 10mph due to a major accident. if the total time of travel was 8 hours,how many miles did he drive at the reduced speed?

OpenStudy (anonymous):

40

OpenStudy (anonymous):

\[t=\frac{d}{r} \]Let s be the distance driven at reduced speed. Solve the following for s:\[\frac{s}{20}+\frac{280-s}{60}=8 \]\[s=100\text{ miles} \]

OpenStudy (anonymous):

robto you have a typo, the slow speed is 10, not 20

OpenStudy (anonymous):

Yup. A blunder of the first order. Thank you for pointing this out.